OpenStudy (anonymous):

What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

\[a\frac{(1-r^{n+1})}{(1-r)}\]

OpenStudy (anonymous):

so would it be 87384?

OpenStudy (anonymous):

\[\sum _{n=0}^{15} (8)(-2)^n\] \[8\sum _{n=0}^{14} (-2)^n\] 8((1-(-2)^15))/(1+2) http://www.wolframalpha.com/input/?i=8%28%281-%28-2%29%5E15%29%29%2F%281%2B2%29 yes

OpenStudy (anonymous):

wanna do some more for me? :)

OpenStudy (anonymous):

no, I think you can handle it

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

yes, you got the right answer

OpenStudy (anonymous):

i googled it... i wanted to see if it was correct

OpenStudy (anonymous):

do some more , you will get the hang of it

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

and if you get stuck , post it ; you will be helped